# Transparent Image Layer Diffusion using Latent Transparency

### v3

@@ 4.8 Raw RGBA Channels
4.9 Community Models
Figure 14: Applying to Community Models. We show that our model can be applied to community LoRAs/Models/PromptStyles to achieve diverse results. All images are achieved using prompt “person”, excepting Animagine using “1girl, masterpiece, fantastic art”.
As shown in Fig. 14, our method can be applied to various community models, LoRAs, and prompt styles, without additional training. More specifically, we try a Minecraft LoRA, a pixel art LoRA, an anime model [8], and several community prompt styles. We can see that applying to different models neither degrades the quality of target model/LoRAs nor degrades the quality of image transparency. This integration capability suggests the model potential for wider use in diverse creative and professional domains.
4.10 Limitations
Figure 15: Limitation. The prompt in this example is “glass cup on table in a warm room”. If the input foreground is a clean transparent object without any illumination or shadow effects, harmonious blending is very difficult since the alpha blending does not create deformation of light or casting of shadow. This can be resolved to some extent when using the background as a condition to generate the foreground. But in this case, getting a clean and reusable transparent object without the influence of illumination is difficult.
@@ Appendix A Padded RGB Channels
$G(\bm{I}_{c})_{p}=\left\{\begin{array}[]{rcl}\phi(\bm{I}_{c})_{p}&&{\text{, if}(\bm{I}_{a})_{p}>0}\\ (\bm{I}_{c})_{p}&&{\text{, otherwise}}\\ \end{array}\right.$ (10)
@@ Appendix B Neural Network Architecture
The latent transparency encoder has exactly same neural network architecture with Stable Diffusion latent VAE encoder [32] (but the input contains 4 channels for RGBA). This model is trained from scratch. The output convolution layer is zero-initialized to avoid initial harmful noise.
@@ Appendix C PatchGAN Discriminator
We use exactly same PatchGAN Discriminator architecture, learning objective, and training scheduling with Latent Diffusion VAE [35]. We directly use the python class LPIPSWithDiscriminator from their official code base (the input channel is set to 4). The generator side objective (from [35]) can be written as
$\mathbb{L}_{\text{disc}}(\bm{z})=\text{relu}(1-D_{\text{disc}}(\bm{z}))\,,$ (11)
@@ Appendix G Background-Conditioned Foregrounds
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 16: Single Transparent Image Results #1. The prompt is “apple”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 17: Single Transparent Image Results #2. The prompt is “a cat”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 18: Single Transparent Image Results #3. The prompt is “a man”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 19: Single Transparent Image Results #4. The prompt is “a man with messy hair”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 20: Single Transparent Image Results #5. The prompt is “woman”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 21: Single Transparent Image Results #6. The prompt is “woman with messy hair”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 22: Single Transparent Image Results #7. The prompt is “dog”. Resolution is $1024\times 1024$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 23: Single Transparent Image Results #8. The prompt is “glass cup”. Resolution is $1024\times 1024$.
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
@@
Figure 24: Single Transparent Image Results #9. The prompt is “dragon”. Resolution is $1152\times 896$.
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
@@
Figure 25: Single Transparent Image Results #10. The prompt is “car”. Resolution is $1152\times 896$.
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
@@
Figure 26: Single Transparent Image Results #11. The prompt is “magic book”. Resolution is $1152\times 896$.
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
@@
Figure 27: Single Transparent Image Results #12. The prompt is “shark”. Resolution is $1024\times 1024$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 28: Single Transparent Image Results #13. The prompt is “magic stone”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 29: Single Transparent Image Results #14. The prompt is “parrot, green fur”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 30: Single Transparent Image Results #15. The prompt is “cyber steampunk robot”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 31: Single Transparent Image Results #16. The prompt is “necromancer”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 32: Multi-layer Results #1. The prompts are “plant on table”, “woman in room”, “dog on floor”, “man walking on street”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 33: Multi-layer Results #2. The prompts are “dog in garden”, “man in street”, “woman, closeup”, “plants on table”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 34: Multi-layer Results #3. The prompts are “cat on floor”, “woman in room”, “man in room”, “golden cup”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 35: Foreground-conditioned Background Results #1. The left-most images are inputs. The prompts are “man sitting on chair”, “man sitting in forest”, “pots on wood table”, “parrot in room”, “parrot in forest”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 36: Foreground-conditioned Background Results #2. The left-most images are inputs. The prompts are “magic book of death”, “magic book of life”, “blue and white porcelain vase in my home”, “blue and white porcelain vase in the museum”, “the man in the snow”, “god of infinity”. Resolution is $896\times 1152$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 37: Background-conditioned Foreground Results #1. The left-most images are inputs. The prompts are “woman climbing mountain”, “man climbing mountain”, “robot in sofa waving hand”, “man in sofa”, “bird on hand”, “apple on hand”. Resolution is $896\times 1152$.

### v4

@@ 4.8. Raw RGBA Channels
4.9. Robust Decoder with Data Augmentations
Fig. 15 shows that we can use data augmentation methods to achieve more robust decoder to handle situations when the latent offset is missing or wrong. This can be useful when certain community models (e.g., anime, cartoon, etc.) fail to produce the desired latent offset during the diffusion process, and we still want to decode useful transparent images from those slightly mismatched latent spaces yielded by those fine-tuned models. To be specific, Fig. 15 simply dropout 30% offsets when training the decoder.
4.10. Community Models
Figure 16. Applying to Community Models. We show that our model can be applied to community LoRAs/Models/PromptStyles to achieve diverse results. All images are achieved using prompt “person”, excepting Animagine using “1girl, masterpiece, fantastic art”.
As shown in Fig. 16, our method can be applied to various community models, LoRAs, and prompt styles, without additional training. More specifically, we try a Minecraft LoRA, a pixel art LoRA, an anime model (cagliostrolab, 2024), and several community prompt styles. We can see that applying to different models neither degrades the quality of target model/LoRAs nor degrades the quality of image transparency. This integration capability suggests the model potential for wider use in diverse creative and professional domains.
4.11. Inference Speed
Table 3. Inference Speed. We report on the inference speed of our framework under different base diffusion models and architectures. We test with a Nvidia RTX 3070 device and a RTX 4090 device. All models use 30 diffusion sampling steps. The reported data are averages of 64 runs. All two-stage model needs to be run twice for two layers, doubling the inference time. This speed may be affected by different inference software.
Candidate SD 1.5 SDXL
Only transparent image (RTX4090) 1.85s 7.3s
Only transparent image (RTX3070) 4.13s 12.5s
Two layers (joint, RTX4090) 5.71s 21.5s
Two layers (joint, RTX3070) 13.01s 41.28s
Two layers (two-stage method, RTX4090) 1.92s $\times$ 2 4.37s $\times$ 2
Two layers (two-stage method, RTX3070) 4.77s $\times$ 2 14.71s $\times$ 2
In Table 3, we report the inference speed with different base diffusion models and architectures. All tests are based on personal level computation devices. We tested SD1.5 and SDXL with Nvidia RTX 3070 and RTX 4090. Our tests include generating single transparent images, generating multiple layers jointly, and generating multiple layers using the two-stage pipelines.
4.12. Limitations
Figure 17. Limitation. The prompt in this example is “glass cup on table in a warm room”. If the input foreground is a clean transparent object without any illumination or shadow effects, harmonious blending is very difficult since the alpha blending does not create deformation of light or casting of shadows. This can be resolved to some extent when using the background as a condition to generate the foreground. But in this case, getting a clean and reusable transparent object without the influence of illumination is difficult.
@@ 5. Conclusion
Acknowledgements.
This work was partially supported by Google through their affiliation with Stanford Institute for Human-centered Artificial Intelligence (HAI).
@@ Appendix A Padded RGB Channels
(10) $G(\bm{I}_{c})_{p}=\left\{\begin{array}[]{rcl}\phi(\bm{I}_{c})_{p}&&{\text{, if}(\bm{I}_{a})_{p}=0}\\ (\bm{I}_{c})_{p}&&{\text{, otherwise}}\\ \end{array}\right.$
@@ Appendix B Neural Network Architecture
The latent transparency encoder has exactly same neural network architecture with Stable Diffusion latent VAE encoder (Podell et al., 2023) (but the input contains 4 channels for RGBA). This model is trained from scratch. The output convolution layer is zero-initialized to avoid initial harmful noise.
@@ Appendix C PatchGAN Discriminator
We use exactly same PatchGAN Discriminator architecture, learning objective, and training scheduling with Latent Diffusion VAE (Rombach et al., 2022). We directly use the python class LPIPSWithDiscriminator from their official code base (the input channel is set to 4). The generator side objective (from (Rombach et al., 2022)) can be written as
(11) $\mathbb{L}_{\text{disc}}(\bm{z})=\text{relu}(1-D_{\text{disc}}(\bm{z}))\,,$
@@ Appendix G Background-Conditioned Foregrounds
Figure 18. Single Transparent Image Results #1. The prompt is “apple”. Resolution is $896\times 1152$.
@@
Figure 19. Single Transparent Image Results #2. The prompt is “a cat”. Resolution is $896\times 1152$.
@@
Figure 20. Single Transparent Image Results #3. The prompt is “a man”. Resolution is $896\times 1152$.
@@
Figure 21. Single Transparent Image Results #4. The prompt is “a man with messy hair”. Resolution is $896\times 1152$.
@@
Figure 22. Single Transparent Image Results #5. The prompt is “woman”. Resolution is $896\times 1152$.
@@
Figure 23. Single Transparent Image Results #6. The prompt is “woman with messy hair”. Resolution is $896\times 1152$.
@@
Figure 24. Single Transparent Image Results #7. The prompt is “dog”. Resolution is $1024\times 1024$.
@@
Figure 25. Single Transparent Image Results #8. The prompt is “glass cup”. Resolution is $1024\times 1024$.
@@
Figure 26. Single Transparent Image Results #9. The prompt is “dragon”. Resolution is $1152\times 896$.
@@
Figure 27. Single Transparent Image Results #10. The prompt is “car”. Resolution is $1152\times 896$.
@@
Figure 28. Single Transparent Image Results #11. The prompt is “magic book”. Resolution is $1152\times 896$.
@@
Figure 29. Single Transparent Image Results #12. The prompt is “shark”. Resolution is $1024\times 1024$.
@@
Figure 30. Single Transparent Image Results #13. The prompt is “magic stone”. Resolution is $896\times 1152$.
@@
Figure 31. Single Transparent Image Results #14. The prompt is “parrot, green fur”. Resolution is $896\times 1152$.
@@
Figure 32. Single Transparent Image Results #15. The prompt is “cyber steampunk robot”. Resolution is $896\times 1152$.
@@
Figure 33. Single Transparent Image Results #16. The prompt is “necromancer”. Resolution is $896\times 1152$.
@@
Figure 34. Multi-layer Results #1. The prompts are “plant on table”, “woman in room”, “dog on floor”, “man walking on street”. Resolution is $896\times 1152$.
@@
Figure 35. Multi-layer Results #2. The prompts are “dog in garden”, “man in street”, “woman, closeup”, “plants on table”. Resolution is $896\times 1152$.
@@
Figure 36. Multi-layer Results #3. The prompts are “cat on floor”, “woman in room”, “man in room”, “golden cup”. Resolution is $896\times 1152$.
@@
Figure 37. Foreground-conditioned Background Results #1. The left-most images are inputs. The prompts are “man sitting on chair”, “man sitting in forest”, “pots on wood table”, “parrot in room”, “parrot in forest”. Resolution is $896\times 1152$.
@@
Figure 38. Foreground-conditioned Background Results #2. The left-most images are inputs. The prompts are “magic book of death”, “magic book of life”, “blue and white porcelain vase in my home”, “blue and white porcelain vase in the museum”, “the man in the snow”, “god of infinity”. Resolution is $896\times 1152$.
@@
Figure 39. Background-conditioned Foreground Results #1. The left-most images are inputs. The prompts are “woman climbing mountain”, “man climbing mountain”, “robot in sofa waving hand”, “man in sofa”, “bird on hand”, “apple on hand”. Resolution is $896\times 1152$.